What is the percent volume for a solution where 175 mL of methanol is added to water to produce 753 mL of solution?
1
Expert's answer
2020-11-17T10:41:10-0500

%(v/v) = "\frac{volume of solute}{total volume of solution}" × 100


which is equal to "\frac{175}{753}" ×100 = 23.24%